# Med-R1: Reinforcement Learning for Generalizable Medical Reasoning in Vision-Language Models

### v2

@@ Abstract
Vision-language models (VLMs) have made significant progress in reasoning within natural scenes, yet their potential in medical imaging remains largely underexplored. Medical reasoning tasks, which require robust image analysis and the generation of well-justified answers, present unique challenges due to the inherent complexity of medical images. Transparency and trustworthiness are critical not only for gaining clinicians’ confidence but also for meeting stringent regulatory requirements. To address these challenges, we propose Med-R1, a novel framework that investigates whether reinforcement learning (RL) can enhance the generalizability and trustworthiness of VLMs in medical reasoning. Building on the recently introduced DeepSeek strategy, we adopt Group Relative Policy Optimization (GRPO) for RL, which encourages models to explore reasoning paths guided by reward signals. Unlike supervised fine-tuning (SFT), which often overfits to training data and struggles with generalization, RL enables models to develop more robust and diverse reasoning capabilities. We comprehensively evaluate Med-R1 across eight distinct medical imaging modalities: Computed Tomography (CT), Magnetic Resonance Imaging (MRI), Ultrasound, Dermoscopy, Fundus Photography, Optical Coherence Tomography (OCT), Microscopy Images, and X-ray Imaging. Compared to the base model, Qwen2-VL-2B, Med-R1 achieves a 29.94% improvement in average accuracy across these modalities and even outperforms Qwen2-VL-72B—a model with 36 times more parameters. To assess model’s generalization abilities, we further test on five different question types: modality recognition, anatomy identification, disease diagnosis, lesion grading, and biological attribute analysis. Med-R1 also demonstrates superior cross-task generalization, outperforming Qwen2-VL-2B by 32.06% and Qwen2-VL-72B in question-type generalization accuracy. These results highlight that RL not only enhances medical reasoning capabilities but also enables parameter-efficient models to exceed the performance of significantly larger counterparts. Combined with its interpretable reasoning outputs, Med-R1 establishes a new potential direction for developing generalizable, trustworthy, and clinically deployable medical VLMs.
@@ 1 Introduction
Vision-language models (VLMs) have achieved remarkable progress in reasoning over natural images (e.g., GPT-4o [16], Gemini-1.5 [31], and Qwen-VL [2]), demonstrating impressive capabilities in tasks such as visual question answering (VQA) and multimodal dialogue. These advancements have been driven by large-scale vision-language pretraining and supervised fine-tuning (SFT), which enable models to associate visual content with textual information. However, applying VLMs to medical imaging reasoning remains a significant challenge. Unlike natural images, medical data demand precise interpretation, requiring VLMs to generate not only final conclusions but also step-by-step reasoning processes that align with clinical logic. For instance, diagnosing a lung nodule in a CT scan necessitates localizing lesions, analyzing morphological features, and integrating patient history—a reasoning chain that must be both accurate and interpretable to gain clinicians’ trust. Furthermore, the diversity of medical imaging modalities (e.g., CT, MRI, microscopy) and task types (e.g., diagnosis, lesion grading) poses stringent demands on model generalizability. These challenges raise a critical question: How can we equip VLMs with robust reasoning abilities that generalize across medical domains while maintaining clinical trustworthiness?
@@
In this paper, we identify that the limitations of current medical VLMs primarily stem from the inherent drawbacks of Supervised Fine-Tuning (SFT) [6, 1]. While SFT has been widely adopted to adapt foundation models to medical imaging [23, 34, 18, 5, 21], it suffers from two fundamental issues that hinder clinical applicability. First, SFT inherently biases models toward memorizing task-specific shortcuts rather than learning generalizable reasoning. By directly aligning model outputs with final answers (e.g., diagnostic labels), SFT encourages overfitting to superficial patterns in training data. Second, the scarcity of high-quality Chain-of-Thought (CoT) annotations severely limits the effectiveness of SFT in medical reasoning. Unlike general-domain tasks, where large-scale CoT datasets can be crowdsourced, medical reasoning requires domain-specific logical structuring (e.g., systematically ruling out differential diagnoses before confirming malignancy). However, curating such CoT datasets is prohibitively expensive, as it demands meticulous annotation by experienced medical professionals to ensure diagnostic validity and clinical coherence. As a result, existing SFT-based medical VLMs [23, 34] lack access to high-quality CoT data, leading to shallow and non-clinically grounded reasoning. These models frequently produce "black-box" predictions, failing to generate interpretable rationales or adapt to out-of-domain tasks (e.g., from radiology to pathology). This lack of transparency and robustness poses a significant challenge to clinical adoption, where explainability and reliability are paramount.
@@
To address these challenges, we propose Med-R1, a framework leveraging Reinforcement Learning (RL) [27] to enhance both the generalizability and trustworthiness of medical VLMs. Unlike SFT, RL guides models to explore diverse reasoning pathways through reward signals, encouraging them to develop intrinsic logical steps rather than relying on annotated CoT rationales. This approach mitigates overfitting and shortcut learning, as demonstrated by Chu et al. [9], who showed that RL-trained vision-language models significantly outperform SFT in cross-domain generalization. Building on recent advancements in RL optimization, we adopt Group Relative Policy Optimization (GRPO) [28], a novel strategy that eliminates the need for complex value models required by traditional methods like Proximal Policy Optimization (PPO) [27]. GRPO incorporates rule-based rewards derived from medical guidelines (e.g., diagnostic criteria for tumors) and group-relative comparisons to stabilize policy updates. For example, rewards can be designed to prioritize reasoning steps that align with radiological decision trees, ensuring clinically plausible rationales. This dual mechanism not only reduces computational overhead but also enforces adherence to domain-specific constraints, making GRPO uniquely suited for medical reasoning tasks where transparency and scalability are paramount.
@@
Figure 1: Overview of the Evaluation Framework: Eight Medical Imaging Modalities and Five Medical Vision Question Answering Tasks. We evaluate Med-R1 across eight distinct medical imaging modalities—CT, MRI, Ultrasound, Dermoscopy, Fundus Photography, OCT, Microscopy Images, and X-ray Imaging—as well as five medical vision question answering tasks: anatomy identification, disease diagnosis, lesion grading, modality recognition, and biological attribute analysis. Example images and corresponding clinical questions illustrate the diversity of medical data and the challenges in developing generalizable vision-language models for automated medical reasoning.
@@
Med-R1 is a generalizable medical VLM that generates interpretable reasoning explanations for medical VQA across eight imaging modalities (Figure 2). We summarized our contribution as follows:
@@
    Emergent Reasoning via Clinically Constrained RL: We adopt Group Relative Policy Optimization, a RL strategy that eliminates the need for explicit reasoning supervision. GRPO integrates rule-based rewards (e.g., enforcing adherence to diagnostic guidelines) and group-relative comparisons, enabling models to learn clinically plausible reasoning paths from final-answer labels alone.
@@
    Robust Generalization with Efficiency: Med-R1 outperforms foundation model by 29.94% and SFT baselines by 15.84% in cross-modality accuracy. We also enhance 32.06% on foundation modal and 11.25% than SFT baselines in cross-task generalization. Moreover, Med-R1 surpasses other larger models in the general domain or medical domain including Qwen2-VL-72B, and MedVInT-7B, demonstrating scalability and trustworthiness for real-world deployment.
@@ 2 Related Works
General VLMs and Medical VLMs. VLMs have significantly advanced multimodal understanding by aligning visual and textual representations through large-scale pretraining. General-purpose VLMs, such as CLIP [26] and BLIP-2 [19], excel in natural image-text tasks but encounter challenges when applied to domain-specific medical reasoning. Recent efforts have sought to adapt these models to medical imaging through SFT on domain-specific datasets. For instance, LLaVA [22] introduces instruction tuning for general-purpose vision-language tasks, achieving strong performance on natural image benchmarks. Building on this, LLaVA-Med [18] fine-tunes LLaVA using radiology reports, demonstrating improved performance on medical VQA tasks. Similarly, Med-Flamingo [23] extends Flamingo by incorporating domain-specific pretraining for medical VQA. However, these SFT-based approaches often suffer from overfitting to narrow medical corpora, limiting their generalizability across diverse imaging modalities and clinical tasks. Our work addresses this gap by introducing RL for generalization, enabling robust performance without modality-specific retraining.
Reinforcement Learning for Post-Training. Reinforcement learning [30, 8] has emerged as a powerful tool for aligning large language models (LLMs) with human preferences through reward signals. Approaches like RLHF [24] and RLAIF [4] leverage human or AI-generated feedback to refine model behavior. In the vision-language domain, RL has been applied to improve VQA accuracy [9] and reduce hallucination [28]. However, existing methods rely on complex reward models or extensive human annotation, limiting scalability in data-scarce domains like medicine. The Group Relative Policy Optimization addresses these limitations by incorporating rule-based rewards derived from clinical guidelines and group-relative comparisons, enabling parameter-efficient adaptation to diverse medical tasks.
Medical Reasoning and Interpretability Recent work emphasizes the importance of interpretable reasoning in medical AI, with approaches like CoT prompting [33] and program-guided reasoning [6]. However, acquiring such annotations in the medical domain is particularly challenging due to the need for expert involvement [20], high annotation costs, and the complexity of medical reasoning. The RL approach enables emergent reasoning without explicit supervision, addressing this challenge.
@@ 4.2 Cross-Modality Generalization
Table 3: Cross-Task Generalization of Med-R1: Performance is evaluated across five clinical reasoning task types (rows: training tasks, columns: test tasks), with darker cell shading indicating stronger generalization. Results demonstrate that domain-specific training (e.g., Disease Diagnosis) preserves in-task expertise while maintaining adaptability to unseen tasks, particularly for modality-agnostic skills like Modality Recognition.
@@ 4.3 Cross-Task Generalization
We also evaluate Med-R1’s generalization across 5 important clinical tasks [14]: Anatomy Identification, Disease Diagnosis, Lesion Grading, Modality Recognition, and Other Attributes. Similar to subsection 4.2, we focus our evaluation on two aspects: cross-modality generalization and comparison against SFT and zero-shot with other VLMs.
@@
Table 4: Performance Comparison of VLMs on Five Medical VQA Tasks: GRPO Fine-Tuning Outperforms Zero-Shot and SFT Baselines Across Diverse Reasoning Tasks. Performance is evaluated on five clinical reasoning types (columns) across three model categories: general-purpose VLMs (zero-shot), medical VLMs (zero-shot), and fine-tuned VLMs (SFT vs. GRPO). Results demonstrate that our compact GRPO-tuned 2B model surpasses even 36× larger zero-shot models and achieves stronger cross-task generalization compared to SFT. Meanwhile, in each column, the best and second-best performances are marked in red and blue.
@@ 5 Conclusion
In this work, we introduce Med-R1, a RL-enhanced VLM designed to improve medical reasoning across diverse imaging modalities. By leveraging GRPO, Med-R1 effectively overcomes the limitations of SFT, significantly enhancing both cross-modality and cross-task generalization. Comprehensive evaluations across eight medical imaging modalities and five clinical reasoning tasks demonstrate that Med-R1 consistently outperforms baseline models, including zero-shot general-purpose VLMs, specialized medical VLMs, and even larger-scale models.
@@
While VLM medical reasoning remains an emerging field with ongoing challenges, Med-R1 highlights the potential of reinforcement learning in developing safer, more transparent, and clinically reliable AI-driven healthcare solutions. Future research could further enhance Med-R1 by integrating additional multi-modal clinical data, improving uncertainty quantification, and refining expert-aligned interpretability, ultimately advancing scalable and trustworthy medical reasoning systems.

### v3

@@ Abstract
Vision-language models (VLMs) have made significant progress in reasoning within natural scenes, yet their potential in medical imaging remains largely underexplored. Medical reasoning tasks, which require robust image analysis and the generation of well-justified answers, present unique challenges due to the inherent complexity of medical images. To be clinically useful, model outputs must be accurate, and logically consistent, as errors or inconsistencies can have critical consequences in medical applications. To address these challenges, we propose Med-R1, a novel framework that investigates whether reinforcement learning (RL) can enhance the generalizability and trustworthiness of VLMs in medical reasoning. Building on the recently introduced DeepSeek strategy, we adopt Group Relative Policy Optimization (GRPO) for RL, which encourages models to explore reasoning paths guided by reward signals. Unlike supervised fine-tuning (SFT), which often overfits to training data and struggles with generalization, RL enables models to develop more robust and diverse reasoning capabilities. We comprehensively evaluate Med-R1 across eight distinct medical imaging modalities: Computed Tomography (CT), Magnetic Resonance Imaging (MRI), Ultrasound, Dermoscopy, Fundus Photography, Optical Coherence Tomography (OCT), Microscopy Images, and X-ray Imaging. Compared to the base model, Qwen2-VL-2B, Med-R1 achieves a 29.94% improvement in average accuracy across these modalities and even outperforms Qwen2-VL-72B—a model with 36 times more parameters. To assess model’s generalization abilities, we further test on five different question types: modality recognition, anatomy identification, disease diagnosis, lesion grading, and biological attribute analysis. Med-R1 also demonstrates superior cross-task generalization, outperforming Qwen2-VL-2B by 32.06% and Qwen2-VL-72B in question-type generalization accuracy. These results highlight that RL not only enhances medical reasoning capabilities but also enables parameter-efficient models to exceed the performance of significantly larger counterparts. By demonstrating strong cross-domain and cross-tasks performance, Med-R1 points toward a new direction for developing practical and generalizable medical VLMs.
@@ 1 Introduction
Vision-language models (VLMs) have achieved remarkable progress in reasoning over natural images (e.g., GPT-4o [16], Gemini-1.5 [33], and Qwen-VL [2]), demonstrating impressive capabilities in tasks such as visual question answering (VQA) and multimodal dialogue. These advancements have been driven by large-scale vision-language pretraining and supervised fine-tuning (SFT), which enables models to associate visual content with textual information. However, applying VLMs to medical imaging reasoning remains a significant challenge. Unlike natural images, medical data demands precise interpretation, requiring VLMs to produce not only answer outcomes but also intermediate reasoning paths that are consistent with clinical decision-making logic. For instance, diagnosing a lung nodule in a CT scan necessitates localizing lesions, analyzing morphological features, and integrating patient history—a reasoning chain that must be both accurate and interpretable to gain clinicians’ trust. Furthermore, the diversity of medical imaging modalities (e.g., CT, MRI, microscopy) and task types (e.g., diagnosis, lesion grading) poses stringent demands on model generalizability. These challenges raise a critical question: How can we equip VLMs with robust reasoning abilities that generalize across medical domains while maintaining medical reliability?
@@
In this paper, we identify that the limitations of current medical VLMs primarily stem from the inherent drawbacks of Supervised Fine-Tuning (SFT) [6, 1]. While SFT has been widely adopted to adapt foundation models to medical imaging [24, 36, 18, 5, 22], it suffers from two fundamental issues that hinder medical applicability. First, SFT inherently biases models toward memorizing task-specific shortcuts rather than learning generalizable reasoning. By directly aligning model outputs with final answers (e.g., diagnostic labels), SFT encourages overfitting to superficial patterns in training data. Second, the scarcity of high-quality Chain-of-Thought (CoT) annotations severely limits the effectiveness of SFT in medical reasoning. Unlike general-domain tasks, where large-scale CoT datasets can be crowdsourced, medical reasoning requires domain-specific logical structuring (e.g., systematically ruling out differential diagnoses before confirming malignancy). However, curating such CoT datasets is prohibitively expensive, as it demands meticulous annotation by experienced medical professionals to ensure diagnostic validity and clinical coherence. As a result, existing SFT-based medical VLMs [24, 36] lack access to high-quality CoT data, leading to shallow and non-clinically grounded reasoning. These models frequently produce "black-box" predictions, struggling to provide traceable reasoning or maintain performance in out-of-domain tasks. This lack of transparency and robustness poses a significant challenge to medical adoption, where explainability and reliability are are indispensable requirements.
@@
To address these challenges, we propose Med-R1, a framework leveraging Reinforcement Learning (RL) [29] to enhance both the generalizability and interpretability of medical VLMs. In contrast to SFT, which primarily aligns outputs with fixed supervision, RL promotes exploration of alternative reasoning strategies by leveraging reward-driven feedback, allowing models to construct internal logic without requiring explicit CoT annotations. This approach mitigates overfitting and shortcut learning, as demonstrated by Chu et al. [9], who showed that RL-trained vision-language models significantly outperform SFT in cross-domain generalization. Building on recent advancements in RL optimization, we adopt Group Relative Policy Optimization (GRPO) [30], a novel strategy that eliminates the need for complex value models required by traditional methods like Proximal Policy Optimization (PPO) [29]. GRPO incorporates rule-based rewards derived from medical guidelines (e.g., diagnostic criteria for tumors) and group-relative comparisons to stabilize policy updates. For example, rewards can be designed to prioritize reasoning steps that align with radiological decision trees, ensuring clinically plausible rationales. This dual mechanism not only reduces computational overhead but also enforces adherence to domain-specific constraints, making GRPO uniquely suited for medical reasoning tasks where interpretability and scalability are essential for deployment.
@@
Figure 1: Overview of the Evaluation Framework: Eight Medical Imaging Modalities and Five Medical Vision Question Answering Tasks. We evaluate Med-R1 across eight distinct medical imaging modalities—CT, MRI, Ultrasound, Dermoscopy, Fundus Photography, OCT, Microscopy Images, and X-ray Imaging—as well as five medical vision question answering tasks: anatomy identification, disease diagnosis, lesion grading, modality recognition, and biological attribute analysis. Example images and corresponding medical questions illustrate the diversity of medical data and the challenges in developing generalizable vision-language models for automated medical reasoning.
@@
Med-R1 is a generalizable medical VLM that generates interpretable reasoning explanations for medical VQA across eight imaging modalities (Figure 2). We summarized the contribution as follows:
@@
    Reward Guided Reasoning via Constrained RL: We employ Group Relative Policy Optimization, a reinforcement learning method that removes the need for explicit reasoning annotations. GRPO combines rule-informed reward signals (e.g., alignment with diagnostic protocols) with group-relative optimization, allowing models to align output behavior with medically grounded reasoning from final-answer supervision alone.
@@
    Robust and Efficient Generalization: Med-R1 outperforms the base foundation model by 29.94% and SFT baselines by 15.84% in cross-modality performance. It further improves 32.06% over the foundation model and 11.25% over SFT baselines on cross-task generalization. Med-R1 also exceeds much larger models in both general-domain and medical-domain settings—including Qwen2-VL-72B and MedVInT-7B—demonstrating scalable reasoning and medical transferability beyond model size.
@@ 2 Related Works
General VLMs and Medical VLMs. VLMs have significantly advanced multimodal understanding by aligning visual and textual representations through large-scale pretraining. General-purpose VLMs, such as CLIP [28] and BLIP-2 [19], excel in natural image-text tasks but encounter challenges when applied to domain-specific medical reasoning. Recent efforts have sought to adapt these models to medical imaging through SFT on domain-specific datasets. For instance, LLaVA [23] introduces instruction tuning for general-purpose vision-language tasks, achieving strong performance on natural image benchmarks. Building on this, LLaVA-Med [18] fine-tunes LLaVA using radiology reports, demonstrating improved performance on medical VQA tasks. Similarly, Med-Flamingo [24] extends Flamingo by incorporating domain-specific pretraining for medical VQA. However, these SFT-based approaches often suffer from overfitting to narrow medical corpora, limiting their generalizability across diverse imaging modalities and medical tasks. Our work addresses this gap by introducing RL for generalization, enabling robust performance without modality-specific retraining.
Reinforcement Learning for Post-Training. Reinforcement learning [32, 8] has emerged as a powerful tool for aligning large language models (LLMs) with human preferences through reward signals. Approaches like RLHF [25] and RLAIF [4] leverage human or AI-generated feedback to refine model behavior. In the vision-language domain, RL has been applied to improve VQA accuracy [9] and reduce hallucination [30]. However, existing methods rely on complex reward models or extensive human annotation, limiting scalability in data-scarce domains like medicine. The Group Relative Policy Optimization addresses these limitations by incorporating rule-based rewards derived from medical guidelines and group-relative comparisons, enabling parameter-efficient adaptation to diverse medical tasks.
Medical Reasoning and Interpretability Recent work emphasizes the importance of interpretable reasoning in medical AI, with approaches like CoT prompting [35] and program-guided reasoning [6]. However, acquiring such annotations in the medical domain is particularly challenging due to the need for expert involvement [21], high annotation costs, and the complexity of medical reasoning. The RL approach enables emergent reasoning without explicit supervision, addressing this challenge. Concurrent to our work, MedVLM-R1 [26] explores reinforcement learning for radiology VQA using a GRPO-based approach. While promising, the study is conducted on a small-scale dataset (600 training samples) and does not evaluate generalization across tasks or modalities, which makes it challenging to assess the robustness of the proposed reasoning capability and its generalizability. Nonetheless, this work also reflects the growing interest in applying reinforcement learning to medical vision-language models.
@@ 3.1 Group Relative Policy Optimization (GRPO)
3.2 No-Thinking-Med-R1
Previous work [20] found that for the image classification task, compared with normal rule-based RL fine-tuning, excluding format accuracy and letting the model only output the final answer could lead to better performance. However, they only discussed this phenomenon for image classification task. In this paper, we further explore the thinking process of medical VQA and propose the No-Thinking-Med-R1.
@@ 4.2 Cross-Modality Generalization
Table 3: Cross-Task Generalization of Med-R1: Performance is evaluated across five medical reasoning task types (rows: training tasks, columns: test tasks), with darker cell shading indicating stronger generalization. Results demonstrate that domain-specific training (e.g., Disease Diagnosis) preserves in-task expertise while maintaining adaptability to unseen tasks, particularly for modality-agnostic skills like Modality Recognition.
@@ 4.3 Cross-Task Generalization
We also evaluate Med-R1’s generalization across 5 important medical tasks [14]: Anatomy Identification, Disease Diagnosis, Lesion Grading, Modality Recognition, and Other Attributes. Similar to subsection 4.2, we focus our evaluation on two aspects: cross-modality generalization and comparison against SFT and zero-shot with other VLMs.
@@
Table 4: Performance Comparison of VLMs on Five Medical VQA Tasks: GRPO Fine-Tuning Outperforms Zero-Shot and SFT Baselines Across Diverse Reasoning Tasks. Performance is evaluated on five medical reasoning types (columns) across three model categories: general-purpose VLMs (zero-shot), medical VLMs (zero-shot), and fine-tuned VLMs (SFT vs. GRPO). Results demonstrate that our compact GRPO-tuned 2B model surpasses even 36× larger zero-shot models and achieves stronger cross-task generalization compared to SFT. Meanwhile, in each column, the best and second-best performances are marked in red and blue.
@@ 5 Conclusion
In this work, we introduce Med-R1, a RL-enhanced VLM designed to improve medical reasoning across diverse imaging modalities. By leveraging GRPO, Med-R1 effectively overcomes the limitations of SFT, significantly enhancing both cross-modality and cross-task generalization. Comprehensive evaluations across eight medical imaging modalities and five medical reasoning tasks demonstrate that Med-R1 consistently outperforms baseline models, including zero-shot general-purpose VLMs, specialized medical VLMs, and even larger-scale models.
@@
While VLM medical reasoning remains an emerging field with ongoing challenges, Med-R1 highlights the potential of reinforcement learning in developing safer, more transparent, and clinically reliable AI-driven healthcare solutions. Future research could further enhance Med-R1 by integrating additional multi-modal medical data, improving uncertainty quantification, and refining expert-aligned interpretability, ultimately advancing scalable and trustworthy medical reasoning systems.